### v1

# Breaking the Bottleneck with DiffuApriel: High-Throughput Diffusion LMs with Mamba Backbone

Abstract
[…]
1 Introduction
The majority of the state-of-the-art Large Language Models (LLMs) (Brown et al., 2020; Achiam et al., 2023; Chowdhery et al., 2023; Touvron et al., 2023a) are trained with the autoregressive (AR) objective: a Transformer predicts the next token given all previous tokens. While simple and powerful, this paradigm faces two persistent limitations. First, AR decoding is inherently sequential: every generated token depends on the full prefix, and thus inference latency grows linearly with output length. Second, AR Transformers are fundamentally constrained by memory: the attention mechanism incurs quadratic complexity in context length, creating computational and memory bottlenecks during both training and auto-regressive inference (Zhou et al., 2024; Tay et al., 2020). These constraints limit throughput and test-time scaling abilitites of these models, motivating research into alternative architectures.
Diffusion Language Models (DLMs) (Li et al., 2025b) provide a flexible alternative to autoregressive generation. Unlike AR decoders, DLMs iteratively denoise entire corrupted sequences in parallel, enabling non-sequential decoding, partial infilling, and semi-autoregressive generation. This opens the door to faster or more controllable generation, especially when fewer denoising steps suffice for acceptable quality. DLMs hold practical promise for scaling iterative reasoning, whether vertically through multiple sampled traces or horizontally through long, coherent single-trace reasoning, as shown in recent AR models (Radhakrishna et al., 2025a, b; Chen et al., 2025). Because DLMs revisit and refine entire sequences during denoising, they can in principle support self-correction by adjusting intermediate errors using both past and future context.
However, a key bottleneck persists: all current DLMs rely on Transformer backbones, making their iterative denoising at inference highly compute-intensive. Each decoding step redundantly recomputes queries, keys, and values (QKV) for every token and layer, even though key–value (KV) states change only marginally between steps. This exhaustive recomputation introduces severe latency and memory overhead, making inference scale with both sequence length and the number of denoising steps, undermining diffusion’s inherent parallelism. The result is a paradox: diffusion promises flexible text generation, yet its efficiency is constrained by the Transformer backbone. Prior work has aimed to accelerate DLMs via improved objectives (Sahoo et al., 2025; Li et al., 2025a; Liu et al., 2025b; Arriola et al., 2025) or Transformer-compatible KV caching (Wu et al., 2025; Ma et al., 2025; Nguyen-Tri et al., 2025), yet the core architectural limitation remains.
In parallel with advances in diffusion language modeling, state-space models (SSMs) have emerged as a powerful class of sequence mixers with linear-time complexity (Gu et al., 2022; Poli et al., 2023; Fu et al., 2023; Gu & Goel, 2023; Dao & Gu, 2024). Recent studies demonstrate that SSMs and SSM-Transformer hybrids can match or even outperform Transformers on language and multimodal benchmarks while achieving substantially higher inference throughput (Gu & Goel, 2023; Somvanshi et al., 2025; Wang et al., 2025a). Despite their potential, SSMs remain unexplored in DLMs. This raises a key question: can structured recurrence serve as an effective language denoiser while enabling faster inference? Exploring the intersection of efficient SSM backbones and masked discrete diffusion for text is the central motivation of our work.
To this end, we introduce the first family of SSM-based discrete masked diffusion language models. Two observations motivate our approach: diffusion LMs are becoming competitive at scale (Gong et al., 2025; Ye et al., 2025; Labs et al., 2025) yet remain tied to Transformer backbones, and recent advances show that Mamba/SSM architectures enable efficient multi-step diffusion in vision (Hu et al., 2024; Ergasti et al., 2025; Ju & Zhou, 2024; Dang et al., 2024). To our knowledge, no prior work has joined these threads for language. Our study fills this gap by showing that SSMs can serve as drop-in denoisers for masked discrete diffusion, thus substantially improving inference efficiency.
Specifically, we replace the Transformer backbone with (i) fully Mamba-2 (Dao & Gu, 2024) denoisers and (ii) hybrid Mamba-2 + Transformer denoisers that retain full-sequence bidirectionality while improving efficiency. Removing quadratic attention reduces per-step latency without altering the probabilistic semantics of masked diffusion. Moreover, block diffusion methods (Arriola et al., 2025; Wu et al., 2025) are orthogonal to our architectural design and can be naturally integrated: in block-DLM hybrids, local Transformer blocks manage short-range dependencies, while SSM layers propagate information globally, combining local precision with efficient long-range reasoning. We leave this direction for future work.
Our contributions are conceptual and empirical rather than algorithmic:
  • New architectural direction. We propose DiffuApriel, which replaces Transformer denoisers with bidirectional Mamba-2 for discrete masked diffusion language modeling. This demonstrates that iterative denoising does not inherently require attention, establishing linear-time backbones as a compelling alternative for scaling diffusion LMs efficiently.
  • Hybrid design exploration. Beyond using only Mamba-2, we also propose DiffuApriel-H, a hybrid architecture that combines attention and state-space layers. We analyze how localized recurrence and global context interact during the denoising process, while still delivering higher throughput than transformer based DLMs. These results provide new insights into the complementary roles of recurrence and attention in diffusion-based generation.
  • Controlled evaluation across scales. We conduct a systematic comparison between DiffuApriel and Transformer-based diffusion models (DiffuTran) under identical training data, tokenization, noise schedules, and decoding steps at three parameter budgets (240M, 0.5B, and 1.3B). Our experiments demonstrate that DiffuApriel and DiffuApriel-H not only improve modeling quality but also deliver substantially higher inference throughput and lower latency, paving the way for more efficient design choices in DLMs.
2 Related Work
Diffusion Language Models (DLMs)
Diffusion modeling replaces left-to-right decoding with iterative denoising. While early work focused on continuous domains (Ho et al., 2020; Song & Ermon, 2021; Rombach et al., 2022; Peebles & Xie, 2023), recent research has adapted diffusion to discrete text. Austin et al. (2021) introduced structured transition matrices with absorbing states, and DiffusionBERT (He et al., 2023) employed a masked corruption process paired with a BERT-style denoiser. Subsequent studies have further refined masking strategies, noise schedules, and sampling procedures (Chen et al., 2023; Lou et al., 2023; Varma et al., 2025; Fathi et al., 2025). Masked diffusion approaching autoregressive quality (Sahoo et al., 2024, 2025) enabled scalable models (Ye et al., 2025; Nie et al., 2025) competitive with LLaMA (Touvron et al., 2023b), with further gains from AR adaptation and instruction tuning (Gong et al., 2025; Zhu et al., 2025). To enhance inference efficiency, recent acceleration techniques (Ma et al., 2025; Wu et al., 2025; Liu et al., 2025c; Israel et al., 2025; Wang et al., 2025b) use approximate KV-caching mechanisms to speed up DLM inference. In contrast, DiffuApriel-H achieves its speedups without relying on KV caching, though its attention layers remain compatible with these optimizations.
State-Space Models (SSMs) for Sequence Modeling
SSMs introduce structured recurrences computable via fast convolutions or scans, achieving linear-time scaling with sequence length. Early models like S4 (Gu et al., 2022) demonstrated strong long-context capacity, while Hyena (Poli et al., 2023) and H3 (Fu et al., 2023) explored alternative structured operators for language modeling. Mamba (Gu & Goel, 2023) extended this line by introducing input-conditioned selective state spaces, matching or surpassing Transformers in accuracy with lower latency and memory use. Somvanshi et al. (2025) further survey the breadth of SSM architectures and highlight their efficiency advantages. In autoregressive language modeling, SSMs rival Transformers at comparable scales while offering higher tokens-per-second throughput. More recent works have evolved various linear-complexity language models and hybrids (Wang et al., 2025a; Ostapenko et al., 2025), ranging from vector recurrences (De et al., 2024) to advanced gating mechanisms (Yang et al., 2024). Yet, these gains remain unexplored in discrete diffusion LMs, where all prior models rely on full-attention denoisers, leaving open whether SSM recurrences can effectively replace attention in iterative denoising for text.
Diffusion with Mamba/SSMs in Vision.
In vision, several studies have replaced Transformer-based DiT backbones with state-space models (SSMs), particularly Mamba variants (Gu & Goel, 2023; Ergasti et al., 2025; Dang et al., 2024), to improve diffusion efficiency. The Diffusion Mamba (DiM) family (Teng et al., 2024; Mo, 2025) achieves high-resolution image synthesis with multi-directional scanning and local feature enhancement, yielding notable throughput gains over attention-based models. VM-DDPM (Ju & Zhou, 2024) fuses convolutional locality with SSM-based global modeling to enhance structural fidelity in medical image generation. ZigMa (Hu et al., 2024) introduces zigzag scanning for faster, more memory-efficient diffusion while maintaining competitive quality, and Phung et al. (2024) incorporates wavelet transforms to strengthen local inductive biases, achieving faster convergence and favorable quality–efficiency trade-offs. Beyond diffusion, Zhu et al. (2024) generalizes Mamba as a vision backbone, and recent surveys (Xu et al., 2024; Liu et al., 2025a; Ergasti et al., 2025; Wang et al., 2024; Dang et al., 2024) map its rapid adoption across segmentation, restoration, and dense prediction. Collectively, these works show that replacing attention with Mamba in multi-step image diffusion preserves or improves quality while substantially lowering per-step compute.
3 Method
3.1 Preliminary
Masked Diffusion Models (MDMs), employ a forward noising process where tokens are progressively replaced by a special [MASK] token (Sahoo et al., 2024; Shi et al., 2024). This process is defined by the transition probability
$q_{t|0}(\mathbf{x}_{t}|\mathbf{x}_{0})=\prod_{i=1}^{L}q_{t|0}(x_{t}^{i}|x_{0}^{i})=\prod_{i=1}^{L}\text{Cat}(x_{t}^{i};(1-t)\delta_{x_{0}^{i}}+t\delta_{\text{MASK}}),$ (1)
where $t\in[0,1]$ controls interpolation between the original data $\mathbf{x}_{0}$ (at $t=0$) and a fully masked sequence (at $t=1$). $\text{Cat}(\cdot)$ denotes the categorical distribution. A parametric model $p_{\theta}$ learns the reverse denoising process, and generation starts from all [MASK] and iteratively unmasks by sampling $p_{\theta}(x_{0}^{i}|\mathbf{x}_{t})$.
Recent theory (MDLM (Shi et al., 2024; Sahoo et al., 2024), RADD (Ou et al., 2024)) simplifies training from a variational bound to a reweighted cross-entropy over masked positions
$\mathcal{L}_{\text{MDM}}=\int_{0}^{1}\frac{1}{t}\,\mathbb{E}_{q_{t|0}(\mathbf{x}_{t}|\mathbf{x}_{0})}\left[\sum_{i:x_{t}^{i}=\text{MASK}}-\log p_{\theta}(x_{0}^{i}|\mathbf{x}_{t})\right]dt.$ (2)
This formulation scales to LLMs as DLMs, with LLaDA (Nie et al., 2025) and Dream-7B (Ye et al., 2025) matching autoregressive performance while enabling parallel decoding and flexible infilling.
The analytical reverse of the forward process defined in Equation 1 is computationally inefficient for generation, as it typically modifies only a single token at each step (Campbell et al., 2022; Lou et al., 2023). A common strategy to accelerate sampling is to employ an MCMC-based approximation of the reverse process (Nie et al., 2025), enabling the model to update multiple masked tokens in a single step when transitioning from a noise level $t$ to an earlier level $s<t$. This yields the following factorized form:
$q_{s|t}=\prod_{i=0}^{n-1}q_{s|t}(\boldsymbol{x}_{s}^{i}\mid\boldsymbol{x}_{t}),\quad q_{s|t}(\boldsymbol{x}_{s}^{i}\mid\boldsymbol{x}_{t})=\begin{cases}1,&\boldsymbol{x}_{t}^{i}\neq[\text{MASK}]\ \text{and}\ \boldsymbol{x}_{s}^{i}=\boldsymbol{x}_{t}^{i}\\ \frac{s}{t},&\boldsymbol{x}_{t}^{i}=[\text{MASK}]\ \text{and}\ \boldsymbol{x}_{s}^{i}=[\text{MASK}]\\ \frac{t-s}{t}\,q_{0|t}(\boldsymbol{x}_{s}^{i}\mid\boldsymbol{x}_{t}),&\boldsymbol{x}_{t}^{i}=[\text{MASK}]\ \text{and}\ \boldsymbol{x}_{s}^{i}\neq[\text{MASK}].\end{cases}$ (3)
Here, $q_{0|t}(\boldsymbol{x}_{s}^{i}\mid\boldsymbol{x}_{t})$ denotes the model-provided distribution over the vocabulary for predicting a non-$\text{MASK}]$ token when $\boldsymbol{x}_{t}^{i}$ is masked. In conditional generation settings—e.g., generating a response $\boldsymbol{x}_{0}$ given a prompt pp—the reverse diffusion process in Equation 3 must be adapted. In this case, the model’s predictive distribution for unmasking a token becomes prompt-conditioned: $q_{0|t}(\boldsymbol{x}_{s}^{i}\mid\boldsymbol{x}_{t},p),$ reflecting that token predictions now depend on both the intermediate noised sequence and the conditioning prompt.
3.2 Diffusion Mamba Language Models
We propose DiffuApriel  a masked diffusion language model whose denoiser is built on a bidirectional state-space Mamba (BiMamba) backbone. DiffuApriel preserves the probabilistic structure of masked discrete diffusion while replacing the Transformer encoder with a bidirectional Mamba mixer, enabling linear-time inference and substantially lower memory overhead during multi-step denoising. We refer to transformer based DLM in our work as DiffuTran for simplicity.
Let $\mathbf{x}\in\mathbb{R}^{B\times L\times d}$ denote a batch of token embeddings, where $B$ is the batch size, $L$ the sequence length, and $d$ the hidden dimension. Each block is composed of two independent Mamba layers: one processes the sequence in the forward direction, and the other processes the sequence in the reverse direction
$\mathbf{h}^{\rightarrow}_{i}$ $=A_{f}\ast\mathbf{h}^{\rightarrow}_{i-1}+B_{f}\ast\mathbf{x}_{i},\quad i=1,\dots,L,$ (4)
$\mathbf{h}^{\leftarrow}_{i}$ $=A_{b}\ast\mathbf{h}^{\leftarrow}_{i+1}+B_{b}\ast\mathbf{x}_{i},\quad i=L,\dots,1.$ (5)
Here $A_{f},B_{f},A_{b},B_{b}\in\mathbb{R}^{k\times d}$ are learnable state-transition kernels implemented as 1D causal and anti-causal convolutions or efficient scan operations (Gu & Goel, 2023). The two directional feature streams are fused through simple additive integration
$\text{Mamba}(x_{i})=\mathbf{h}_{i}=\mathbf{h}^{\rightarrow}_{i}+\mathbf{h}^{\leftarrow}_{i},$ (6)
providing a symmetric context representation while maintaining numerical stability. The resulting hidden sequence $\mathbf{H}=(\mathbf{h}_{1},\dots,\mathbf{h}_{L})$ is then normalized and passed through a lightweight feed-forward projection before residual addition.
Each Diffusion Block employs timestep-conditioned adaptive layer normalization (AdaLN) to inject the diffusion noise level into the hidden activations. A small MLP maps the scalar timestep $t$ to a continuous embedding $\tau_{t}=\text{MLP}(t)\in\mathbb{R}^{d_{c}}$, which modulates both the mixer and MLP sublayers:
$\text{AdaLN}(\mathbf{x};\tau_{t})=\gamma_{t}\cdot\frac{\mathbf{x}-\mu(\mathbf{x})}{\sigma(\mathbf{x})}+\beta_{t},\quad(\gamma_{t},\beta_{t})=W_{\text{cond}}\tau_{t}.$ (7)
This conditioning allows the denoiser to adapt its internal recurrence dynamics to varying noise levels across diffusion steps.
Figure 1: Schematic diagram of our proposed DiffuApriel architecture where mixer blocks replaces attention layers with bidirectional Mamba layers. In our experiments, to maintain comparability with DiffuTran, we treat the MLP layer as optional and refer to this variant as DiffuApriel+MLP. For DiffuApriel-H we have interleaved attention layers after every $K$ mamba layers. Attention provides global token interactions while Mamba enables efficient state space sequence modeling, allowing the hybrid denoiser to capture both long-range dependencies and local temporal dynamics with significantly improved efficiency. In our experiments, we fix $K=5$.
Each block then applies Mamba mixing followed by an optional MLP refinement as shown in Figure 1:
$\mathbf{y}_{\text{mixer}}$ $=\text{Mamba}(\text{AdaLN}(\mathbf{x};\tau_{t}))+\mathbf{x},$ (8)
$\mathbf{y}_{\text{mlp}}$ $=\text{MLP}(\text{AdaLN}(\mathbf{y}_{\text{mixer}};\tau_{t}))+\mathbf{y}_{\text{mixer}}.$ (9)
The final output $\mathbf{y}_{\text{mlp}}$ is passed to the next block in the stack. During denoising, given masked input embeddings $\mathbf{x}_{t}$ and timestep embedding $\tau_{t}$, the model predicts token logits via the full diffusion stack:
$\mathbf{z}=f_{\theta}(\mathbf{x}_{t},\tau_{t})=\text{OutputProj}\!\left(\mathrm{Block}_{N}(\cdots\mathrm{Block}_{1}(E(\mathbf{x}_{t}),\tau_{t})\cdots)\right).$ (10)
The conditional distribution over clean tokens is then given by:
$p_{\theta}(\mathbf{x}_{0}|\mathbf{x}_{t},t)=\prod_{i=1}^{L}\text{Cat}(x_{0}^{i};\text{softmax}(z_{i})).$ (11)
Training minimizes the masked diffusion objective (Eq. 2), preserving probabilistic consistency with absorbing-state discrete diffusion. By replacing attention with bidirectional state-space dynamics, DiffuApriel achieves efficient and expressive denoising with linear scaling in both sequence length and memory.
Recent results in large-scale autoregressive modeling show that hybrid architectures (Lieber et al., 2024; Bae et al., 2025; Zuo et al., 2025; Wang et al., 2025a) combining attention and state-space layers can outperform both pure Transformer and pure Mamba models under similar compute budgets, leveraging complementary strengths in local recurrence and global dependencies. Motivated by these findings, we investigate hybrid DLMs by interleaving bidirectional Mamba and Transformer blocks, inserting one attention mixer every $K$ Mamba mixer blocks (with $K=5$ in all experiments). Extending this paradigm from AR to diffusion, we examine whether similar complementarities hold under masked denoising objectives.
Model Params (B) Layers $\mathbf{d_{\text{model}}}$ $\mathbf{d_{\text{mlp}}}$ $\mathbf{d_{\text{head}}}$ $\mathbf{d_{\text{state}}}$ Context Len. #\# Tokens
DiffuTran-0.24B 0.24 24 960 960*4 32 - 1024 25B
DiffuApriel-0.29B 0.29 24 960 960*2 32 128 1024 25B
DiffuApriel-H-0.26B (K=5) 0.26 24 960 960*2 32 128 1024 25B
DiffuTran-0.5B 0.52 36 960 960*4 32 - 1024 50B
DiffuApriel-0.6B 0.67 36 960 960*2 32 128 1024 50B
DiffuApriel-H-0.6B (K=5) 0.61 36 960 960*2 32 128 1024 50B
DiffuTran-1.3B 1.3 24 1920 1920*4 32 - 1024 120B
DiffuApriel-1.6B 1.6 24 1920 1920*2 32 128 1024 120B
DiffuApriel-H-1.5B (K=5) 1.5 24 1920 1920*2 32 128 1024 120B
Table 1: Model configurations for DiffuTran, DiffuApriel and DiffuApriel-H in our experiments. All variants are trained under the same diffusion objective, noise schedule, and tokenization. The only difference lies in the internal mixer architecture. The MLP expansion ratio in DiffuApriel+MLP and DiffuApriel-H+MLP configurations is reduced by half for comparable parameter budgets. In the hybrid architecture, an attention layer is inserted every $K$ layers; we set $K=5$ in all experiments.
4 Experiments
We evaluate three denoiser architectures under an identical diffusion objective, masking process, and noise schedule to ensure a controlled comparison. The first configuration employs an attention-based mixer in every diffusion block, the second replaces these with bidirectional Mamba mixers, and the third adopts a hybrid design that interleaves attention and Mamba blocks following the scheduling principles outlined in Section 3.2. Across all variants, the overall architecture, data flow, and conditioning mechanisms remain identical, with only the internal mixer type varying, as illustrated in Figure 1.
To maintain comparable parameter budgets, the MLP expansion ratio in DiffuApriel+MLP and hybrid configurations is set to half that of the attention-only model, while other hyperparameters, such as the number of blocks, hidden size, etc are kept constant. This design ensures a fair, architecture-isolated evaluation of modeling quality, efficiency, and inference throughput. Our goals are to answer:
  1. 1.
    Throughput and Latency: Does an SSM backbone yield higher tokens/sec and lower per-step latency at matched quality?
  2. 2.
    Scaling: Do the efficiency gains persist or improve with model size?
4.1 Data, Tokenization, and Pretraining Setup
We evaluate our DiffuApriel and DiffuApriel-H models as generative language models by comparing their likelihood, throughput, and latency against DiffuTran at three parameter scales: 240M, 0.5B, and 1.3B. In our experiments, to maintain comparability with DiffuTran, we treat the MLP layer as optional and refer to this variant as DiffuApriel+MLP. All models are trained on the DCLM dataset (Li et al., 2024) using the GPT2 tokenizer (Radford et al., 2019), with a fixed context length of 1,024 tokens. We train all models with a fixed log-linear noise schedule and the masked diffusion objective with decoding steps fixed to $S=128$ for all runs, under approximate Quokka optimal compute budget (Ni et al., 2025) as given in Table 1. For measuring inference throughput, we record the wall-clock model latency and tokens-per-second throughput on a single NVIDIA H100 GPU using bf16 precision and PyTorch backend. Each measurement averages over 20 full diffusion decoding runs (128 denoising steps per run) following 5 warm-up iterations, with timings captured via CUDA Graphs to eliminate launch overheads.
4.2 Inference Throughput and Modeling Quality
Figure 2: Inference throughput and Model’s Latency per forward pass vs. sequence length with a batch size of 1. It can be observed that inference throughput in DiffuApriel scales nearly linearly with $L$, whereas DiffuTran degrades super-linearly due to quadratic attention. At 1.3B scale, DiffuApriel and DiffuApriel+MLP yields 5.3x and 4.4x throughput improvement over DiffuTran respectively. Further hybrid architectures, DiffuApriel-H and DiffuApriel-H+MLP gives 2.8x and 2.6x throughput improvement over DiffuTran.
We observe distinct scaling behaviors as the sequence length increases from L=32L{=}32 to L=65L{=}65K between the attention based diffusion model and our proposed DiffuApriel as shown in Figure 2. For DiffuApriel (we use Mamba-2 training kernel for both training and inference), inference throughput grows and saturates for longer sequences. This trend arises from their $\mathcal{O}(B\,L\,d)$ computational and memory complexity, where total latency $t(L)$ also scales linearly with $L$. Once the GPU reaches its peak memory bandwidth, the model becomes memory bound (Baruah et al., 2025), because the work (total tokens $\propto L$) and the latency ($t(L)\propto L$) scale equally, the resulting tokens-per-second rate ($T(L)\propto L/t(L)$) becomes approximately constant: $T_{\max}\approx M_{p}/\text{bytes per token}$, where $M_{p}$ is peak memory bandwidth per second (bytes/sec).
In contrast, DiffuTran exhibit an initial rise followed by a sharp decline in throughput as sequence length increases, owing to their quadratic $\mathcal{O}(B\,L^{2}\,d)$ cost. The quadratic growth of the $QK^{\top}$ and softmax operations means the total latency $t(L)$ scales super-linearly ($t(L)\!\sim\!\beta BL^{2}$). Since throughput is defined as tokens per second ($T(L)\propto L/t(L)$), the $L$ growth in the numerator is overcome by the $L^{2}$ growth in the denominator, causing throughput to decay roughly as $T(L)\!\propto\!1/L$. This quadratic cost is further exacerbated by unfavorable GEMM (General Matrix–Matrix Multiplication) aspect ratios ($L\gg d_{h};\;d_{h}:\text{head dimension}$) at large $L$.
For hybrid diffusion models interleaving state-space and attention layers, the overall complexity falls between the linear $\mathcal{O}(B\,L\,d)$ scaling of Mamba and the quadratic $\mathcal{O}(B\,L^{2}\,d)$ scaling of attention. With attention inserted every $K$ layers, the effective cost approximates $\mathcal{O}(B\,L\,d+B\,L^{2}\,d/K)$. Thus, the linear term dominates for moderate $L$, yielding near-linear throughput, while the quadratic component contributes mildly at larger $L$. This interleaving achieves a balance, preserving attention’s global context modeling while maintaining the high streaming efficiency of Mamba. These observations align with the theoretical cost model and confirm that Mamba-like architectures maintain stable inference efficiency under long-context regimes, whereas attention models face quadratic degradation beyond a critical sequence length.
4.3 Evaluation
We first measure validation perplexity (PPL) for different configuration of models trained on DCLM dataset as shown in Table 2. Across all model sizes, hybrid SSM–Transformer denoisers consistently outperform their pure-attention counterparts under both Chinchilla and Quokka compute budgets. Particularly, at the 1.3B scale, DiffuTran attains 25.01 and 22.72 PPL, but DiffuApriel-H+MLP achieves the best results with 22.89 and 20.17, delivering roughly a 2% perplexity reduction and an accompanying 2.6× inference speedup Figure 2. At smaller scales (240M and 0.5B), DiffuApriel-H and DiffuApriel-H+MLP remain competitive, often securing the best or second-best perplexities, demonstrating that hybridization provides consistent gains across regimes while enabling more efficient diffusion denoising than attention-only DiffuTran.
Configuration 240M 0.5B 1.3B
Chinchilla Quokka Chinchilla Quokka Chinchilla Quokka
DiffuTran 43.82 32.11 31.03 25.07 25.01 22.72
DiffuApriel 44.60 33.96 30.05 25.99 23.55 21.52
DiffuApriel + MLP 44.09 33.01 30.78 25.46 23.96 21.41
DiffuApriel-H 42.17 33.14 28.73 24.68 23.36 21.29
DiffuApriel-H + MLP 42.67 32.49 29.06 25.14 22.89 20.17
[…]
We next assess models’ ability to generalize zero-shot to held-out datasets. Following Sahoo et al. (2024), we evaluate perplexity upper bounds on Penn Treebank (PTB; Marcus et al., 1993), WikiText (Merity et al., 2017), LM1B (Chelba et al., 2013), Lambada (Paperno et al., 2016), AG News (Zhang et al., 2015), and Pubmed/Arxiv subsets of Scientific Papers (Cohan et al., 2018). All models are trained on DCLM and evaluated without fine-tuning.
Model Size Configuration PTB WikiText LM1B Lambada AG News PubMed ArXiv
240M DiffuTran 153.47 49.06 142.25 49.06 105.06 29.83 43.79
DiffuApriel 203.92 54.66 131.48 48.11 111.16 35.14 34.33
DiffuApriel+MLP 99.04 51.35 169.92 46.66 112.59 33.98 33.20
DiffuApriel-H 160.40 51.76 131.98 46.47 107.49 33.38 32.12
DiffuApriel-H+MLP 147.97 50.96 145.25 46.43 114.20 31.94 30.99
0.5B DiffuTran 112.69 40.76 103.51 39.73 81.49 25.51 25.37
DiffuApriel 127.11 44.16 108.59 40.42 90.80 28.32 28.20
DiffuApriel+MLP 116.29 42.56 117.42 40.31 90.10 27.60 27.79
DiffuApriel-H 118.82 42.16 102.87 39.10 80.46 25.58 25.44
DiffuApriel-H+MLP 110.42 40.25 100.59 39.30 78.38 26.12 25.84
1.3B DiffuTran 101.14 36.63 99.72 36.73 67.53 23.03 23.25
DiffuApriel 112.79 35.69 94.73 37.23 65.08 23.44 23.98
DiffuApriel+MLP 99.04 34.74 102.01 36.04 68.75 22.35 22.98
DiffuApriel-H 103.96 34.86 103.52 40.41 65.34 21.81 22.35
DiffuApriel-H+MLP 96.51 31.92 92.83 34.04 63.22 19.05 20.67
Table 3: Zeroshot Perplexities (PPL $\downarrow$) across Benchmarks for Different Model Sizes and Configurations. Best PPL are highlighted in blue and second best are underlined. We observe that DiffuApriel-H (1.3B) achieves the best overall performance across all configurations, with DiffuApriel as the next strongest model. At smaller scales (240M), however, DiffuApriel delivers performance comparable to its attention-based diffusion counterpart, indicating that the advantages of hybridization become more pronounced at larger model sizes.
At the 240M scale, both DiffuApriel and DiffuApriel-H either improve upon or remain comparable to the transformer-based DiffuTran, demonstrating that replacing attention with state-space modeling does not compromise zero-shot generalization at small scale. Notably, DiffuTran achieves the lowest perplexity on only three datasets at this scale, with DiffuApriel and DiffuApriel-H consistently taking the second-best position—showcasing the strong competitiveness of Mamba-based denoisers even without attention.
At 0.5B, the benefits of hybridization become immediately evident: DiffuApriel-H+MLP outperforms all other variants on five out of seven datasets, establishing a clear performance lead. The advantage becomes even more pronounced at 1.3B, where DiffuApriel-H+MLP surpasses DiffuTran across all datasets, with DiffuApriel ranking second in most cases—indicating that Mamba’s sequence-modeling inductive bias scales more effectively than pure attention for diffusion denoising. Moreover, our results indicate that including the optional MLP layer enhances the performance of both DiffuApriel and DiffuApriel-H.
These results highlight a key trend: while Mamba provides efficient long-range token mixing with linear complexity, combining it with interleaved attention layers in DiffuApriel-H captures complementary global dependencies, enabling stronger generalization—especially at larger scales. The hybrid architecture thus offers the best of both worlds: the efficiency and stability of SSMs and the global contextual reasoning of attention, making DiffuApriel-H the most effective design for large-scale DLMs.
Configuration (1.3B) OBQA HS PIQA LOQA ARC Avg.
DiffuTran 26.61 34.97 60.64 21.86 24.98 33.81
DiffuApriel 32.87 37.13 60.17 30.58 27.65 37.68
DiffuApriel+MLP 32.12 37.74 62.56 29.87 28.33 38.12
DiffuApriel-H 31.65 37.19 61.32 30.65 28.94 37.95
DiffuApriel-H+MLP 33.87 38.02 59.13 32.35 27.83 38.24
OBQA: OpenBookQA, LQA: LogicQA, HS: HellaSwag
Table 4: Zero-shot downstream accuracy ($\uparrow$) on reasoning and commonsense benchmarks (OpenBookQA (Mihaylov et al., 2018), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), LogicQA (Liu et al., 2020), and ARC (Clark et al., 2018)) for 1.3B models. Overall performance at this scale remains modest, as 1.3B models are generally not strong on these challenging evaluations. However, both DiffuApriel and DiffuApriel-H consistently outperform the attention-based diffusion baseline (DiffuTran), with DiffuApriel-H and its MLP-enhanced variant achieving the strongest results. Best scores are highlighted in blue and second-best are underlined.
We also report downstream eval results for 1.3B models in Table 4. As expected, the absolute scores remain modest at this scale, reflecting the difficulty of these reasoning and knowledge-intensive benchmarks at 1.3B scale. Nonetheless, a consistent trend emerges across all tasks: DiffuApriel and DiffuApriel-H clearly outperform DiffuTran by $\approx\boldsymbol{4\%}$ on average, indicating that linear-time state space modeling provides a stronger denoising backbone for diffusion-based language models. Further integrating lightweight MLP adapters and hybridizing with attention layers, as in DiffuApriel-H and DiffuApriel-H+MLP, produces the strongest results, suggesting that a small degree of explicit cross-token interaction complements the Mamba backbone. Even at low model capacity where downstream metrics are weak, Mamba based DLM clearly outperform attention based DiffuTran, demonstrating the effectiveness of state space driven denoising and its superior inference throughput.
5 Conclusion
This paper introduced a new architecture for masked DLMs by replacing Transformer denoisers with state space models. We presented DiffuApriel, to our knowledge the first diffusion LM built on a bidirectional Mamba backbone, and DiffuApriel-H, a hybrid variant that adds interleaved attention layers for complementary global context. Across 240M to 1.3B parameters, both models match or surpass the Transformer baseline DiffuTran while delivering significantly higher throughput and lower latency. Our results connect to a growing body of work exploring efficient linear-time mixers, including SSMs, DeltaNet-style token mixers, and hybrid architectures, underscoring that attention is not strictly required for iterative denoising. A remaining limitation of our approach is that performance advantages diminish on short contexts. Looking forward, masked diffusion with hybrid backbones can accommodate block-diffusion techniques in a complementary way. Exploring this integration is a compelling avenue for future work.
Appendix A Downstream Evaluation Results
We also report downstream eval results for 1.3B models in Table 5. As expected, the absolute scores remain modest at this scale, reflecting the difficulty of these reasoning and knowledge-intensive benchmarks at 1.3B scale. Nonetheless, a consistent trend emerges across all tasks: DiffuApriel and DiffuApriel-H clearly outperform DiffuTran by $3.3$ points on average, indicating that linear-time state space modeling provides a stronger denoising backbone for diffusion-based language models. Further integrating lightweight MLP adapters and hybridizing with attention layers, as in DiffuApriel-H and DiffuApriel-H+MLP, produces the strongest results, suggesting that a small degree of explicit cross-token interaction complements the Mamba backbone. Even at low model capacity where downstream metrics are weak, Mamba based diffusion models clearly outperform attention based variants, demonstrating the effectiveness of state space driven denoising and its superior inference throughput.
Configuration (1.3B) OBQA HS PIQA LOQA ARC MMLU Avg.
DiffuTran 26.61 34.97 60.64 21.86 24.98 25.25 32.38
DiffuApriel 32.87 37.13 60.17 30.58 27.65 25.12 35.58
DiffuApriel+MLP 32.12 37.74 62.56 29.87 28.33 26.91 36.25
DiffuApriel-H 31.65 37.19 61.32 30.65 28.94 25.44 35.86
DiffuApriel-H+MLP 33.87 38.02 59.13 32.35 27.83 26.39 36.26
OBQA: OpenBookQA, LQA: LogicQA, HS: HellaSwag
Table 5: Zero-shot downstream accuracy ($\uparrow$) on reasoning and commonsense benchmarks (OpenBookQA (Mihaylov et al., 2018), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), LogicQA (Liu et al., 2020), ARC-C (Clark et al., 2018), and MMLU (Hendrycks et al., 2021)) for 1.3B models. Overall performance at this scale remains modest, as 1.3B models are generally not strong on these challenging evaluations. However, both DiffuApriel and DiffuApriel-H consistently outperform the attention-based diffusion baseline (DiffuTran), with DiffuApriel-H and its MLP-enhanced variant achieving the strongest results. Best scores are highlighted in blue and second-best are underlined.