### v3

# arXiv:2511.15927v3[cs.LG] 27 Feb 2026

Abstract
[…]
Machine Learning, ICML
1 Introduction
The majority of the state-of-the-art Large Language Models (LLMs) (Brown et al., 2020; Achiam et al., 2023; Chowdhery et al., 2023; Touvron et al., 2023a) rely on full multi-head attention mechanism (MHA) (Bahdanau et al., 2014; Vaswani et al., 2017) and are trained with the autoregressive (AR) objective: a Transformer predicts the next token given all previous tokens. While simple and powerful, this paradigm faces several persistent limitations. First, AR decoding is inherently sequential: every generated token depends on the full prefix, and thus inference latency grows linearly with output length. The AR training paradigm also limits data efficiency: each training example provides supervision for only a single next-token prediction, constraining the amount of learning signal that can be extracted from limited data (Ni et al., 2025a). This limitation becomes increasingly relevant as high-quality data, rather than compute, emerges as a key bottleneck in scaling language models. Furthermore, AR Transformers are constrained by both memory and compute: the KV cache grows linearly with context length, inducing increasing memory pressure, while the attention mechanism incurs quadratic compute cost in the sequence length, limiting long-context inference and training (Zhou et al., 2024; Tay et al., 2020). These constraints limit throughput and test-time scaling efficiency, motivating research into alternative architectures.
Diffusion Language Models (DLMs) (Li et al., 2025) provide a flexible alternative to AR generation. Unlike AR decoders, DLMs iteratively denoise entire corrupted sequences in parallel, enabling non-sequential multi-token generation, partial infilling, and self-correction. In principle, this opens the door to faster generation and self-refined reasoning, especially when fewer denoising steps suffice for acceptable quality. However, a key bottleneck persists: all current DLMs rely on Transformer backbones, making their iterative denoising compute or/and memory intensive pushing their throughput well below the AR competitors especially at long sequences. In standard DLMs, denoising proceeds via repeated full sequence re-encoding using bidirectional attention, as each step conditions on both past and future tokens. Because token states evolve across denoising steps, representations cannot be incrementally reused, leading to per-step costs that scale quadratically with sequence length. While several DLM variants enable KV caching (Wu et al., 2025; Ma et al., 2025; Nguyen-Tri et al., 2025; Fathi et al., 2025), the core architectural limitation remains: as the cache grows with sequence length, it induces increasing memory traffic becoming the key bottleneck at inference. Consequently, while inference cost necessarily scales with the number of denoising steps and the sequence length, Transformer-based DLMs incur additional attention and cache related overheads making per token latency grow with sequence length. The result is a paradox: DLMs promise flexible text generation, yet their efficiency is constrained by increasing memory overhead and periodic cache recomputation induced by the Transformer backbone (Wu et al., 2025).
In parallel with advances in diffusion language modeling, state-space models (SSMs) have emerged as a powerful class of sequence mixers with linear-time complexity (Gu et al., 2022; Poli et al., 2023; Fu et al., 2023; Gu & Goel, 2023; Dao & Gu, 2024). Recent studies demonstrate that SSMs and SSM–Transformer hybrids can match or even outperform Transformers while achieving substantially higher inference throughput (Gu & Goel, 2023; Somvanshi et al., 2025; Wang et al., 2025a). Despite their potential, SSMs remain unexplored in DLMs. This raises a key question: can structured recurrence serve as an effective language denoiser while enabling faster inference? Our central motivation is to explore this intersection of efficient SSM backbones and masked discrete diffusion for language modeling.
To this end, we pre-train standard full-sequence DLMs (Sahoo et al., 2024) whose Transformer backbone is modified in two ways: (i) by replacing all Multi-Head Attention (MHA) mixers with Mamba-2 (Dao & Gu, 2024), and (ii) by replacing only a subset of MHA mixers with Mamba-2 to form a hybrid architecture. Further, we employ bidirectional Mamba-2 mixers, as masked diffusion requires conditioning on both past and future context at each denoising step. By removing quadratic attention, we reduce per-step latency and memory pressure without altering the probabilistic semantics of masked diffusion. As a result, Mamba-based DLMs achieve higher throughput than both Transformer-based DLMs and autoregressive baselines across a range of decoding algorithms, including block diffusion (Arriola et al., 2025; Wu et al., 2025).
The main contributions of this work are:
  • New architectural direction. We propose DiffuMamba, which replaces Transformer denoisers with bidirectional Mamba-2 mixers for discrete masked diffusion language modeling. We further introduce DiffuMamba-H, a sparse hybrid variant that interleaves attention by inserting one Transformer block every five Mamba blocks ($\approx 20\%$ attention), starting with an attention block. Together, these models demonstrate that iterative denoising does not inherently require dense attention, positioning linear-time backbones as a scalable alternative for diffusion language models.
  • Controlled evaluation across scales. We conduct a systematic comparison between DiffuMamba and Transformer-based diffusion models (DiffuTran) under identical training data, tokenization, noise schedules, and decoding steps at three parameter budgets (240M, 0.5B, and 1.3B). Our experiments demonstrate that DiffuMamba and DiffuMamba-H can match the modeling quality of the full transformer.
  • Comprehensive throughput benchmarking. We present an extensive asymptotic and empirical analysis of modern DLM inference strategies, scaling generation length beyond 100k tokens, a regime particularly relevant for complex reasoning workloads. Across all evaluated inference algorithms, Mamba-backed DLMs consistently outperform Transformer-based counterparts: by up to 8.2× in full-sequence denoising (LABEL:fig:inf:a and LABEL:fig:inf:b) and by 2.3× in the most efficient block-wise autoregressive denoising setting (LABEL:fig:infb:a and LABEL:fig:infb:b). Notably, the latter is the only inference strategy that scales linearly with sequence length, and when combined with Mamba-based denoisers, it outperforms all other baselines, outlining a promising path for future diffusion-based generation systems.
2 Related Work
Diffusion Language Models (DLMs)
Diffusion modeling replaces left-to-right decoding with iterative denoising. While early work focused on continuous domains (Ho et al., 2020; Song & Ermon, 2021; Rombach et al., 2022; Peebles & Xie, 2023), recent research has adapted diffusion to discrete text. Austin et al. (2021) introduced structured transition matrices with absorbing states, and DiffusionBERT (He et al., 2023) employed a masked corruption process paired with a BERT-style denoiser. Subsequent studies have further refined masking strategies, noise schedules, and sampling procedures (Chen et al., 2023; Lou et al., 2023; Varma et al., 2025; Fathi et al., 2025). Masked diffusion approaching autoregressive quality (Sahoo et al., 2024, 2025) enabled scalable models (Ye et al., 2025; Nie et al., 2025) competitive with LLaMA (Touvron et al., 2023b), with further gains from AR adaptation and instruction tuning (Gong et al., 2025; Zhu et al., 2025). To enhance inference efficiency, recent acceleration techniques (Ma et al., 2025; Wu et al., 2025; Liu et al., 2025b; Israel et al., 2025; Wang et al., 2025b) use approximate KV-caching mechanisms to speed up DLM inference. In contrast, DiffuMamba-H achieves its speedups without relying on KV caching, though its attention layers remain compatible with these optimizations.
State-Space Models (SSMs) for Sequence Modeling
SSMs introduce structured recurrences computable via fast convolutions or scans, achieving linear-time scaling with sequence length. Early models like S4 (Gu et al., 2022) demonstrated strong long-context capacity, while Hyena (Poli et al., 2023) and H3 (Fu et al., 2023) explored alternative structured operators for language modeling. Mamba (Gu & Goel, 2023) extended this line by introducing input-conditioned selective state spaces, matching or surpassing Transformers in accuracy with lower latency and memory use. Somvanshi et al. (2025) further survey the breadth of SSM architectures and highlight their efficiency advantages. In autoregressive language modeling, SSMs rival Transformers at comparable scales while offering higher tokens-per-second throughput. More recent works have evolved various linear-complexity LMs and hybrids (Wang et al., 2025a; Ostapenko et al., 2025), ranging from vector recurrences (De et al., 2024) to advanced gating mechanisms (Yang et al., 2024). Although SSM backbones have been used in diffusion models for non-textual domains such as generative modelling in biological sequences (Sahoo et al., 2024), text diffusion LMs continue to rely on full-attention denoisers, leaving open whether SSM recurrences can replace attention for iterative denoising.
Diffusion with Mamba/SSMs in Vision.
In vision, several studies have replaced Transformer-based DiT backbones with state-space models (SSMs), particularly Mamba variants (Gu & Goel, 2023; Ergasti et al., 2025; Dang et al., 2024), to improve diffusion efficiency. The Diffusion Mamba (DiM) family (Teng et al., 2024; Mo, 2025) achieves high-resolution image synthesis with multi-directional scanning and local feature enhancement, yielding notable throughput gains over attention-based models. VM-DDPM (Ju & Zhou, 2024) fuses convolutional locality with SSM-based global modeling to enhance structural fidelity in medical image generation. ZigMa (Hu et al., 2024) introduces zigzag scanning for faster, more memory-efficient diffusion while maintaining competitive quality, and Phung et al. (2024) incorporates wavelet transforms to strengthen local inductive biases, achieving faster convergence and favorable quality–efficiency trade-offs. Beyond diffusion, Zhu et al. (2024) generalizes Mamba as a vision backbone, and recent surveys (Xu et al., 2024; Liu et al., 2025a; Ergasti et al., 2025; Wang et al., 2024; Dang et al., 2024) map its rapid adoption across segmentation, restoration, and dense prediction. Collectively, these works show that replacing attention with Mamba in multi-step image diffusion preserves or improves quality while substantially lowering per-step compute.
3 Method
3.1 Preliminary
Masked Diffusion Models (MDMs), employ a forward noising process where tokens are progressively replaced by a special [MASK] token (Sahoo et al., 2024; Shi et al., 2024). This process is defined by the transition probability
$q_{t|0}(\mathbf{x}_{t}\mid\mathbf{x}_{0})=\prod_{i=1}^{L}q_{t|0}(x_{t}^{i}\mid x_{0}^{i})$
$\prod_{i=1}^{L}\text{Cat}\!\left(x_{t}^{i};\;(1-t)\delta_{x_{0}^{i}}+t\,\delta_{\text{MASK}}\right),$ (1)
where $t\in[0,1]$ controls interpolation between the original data $\mathbf{x}_{0}$ (at $t=0$) and a fully masked sequence (at $t=1$). $\text{Cat}(\cdot)$ denotes the categorical distribution. A parametric model $p_{\theta}$ learns the reverse denoising process, and generation starts from all [MASK] and iteratively unmasks by sampling $p_{\theta}(x_{0}^{i}|\mathbf{x}_{t})$.
Recent theory (MDM (Shi et al., 2024; Sahoo et al., 2024), RADD (Ou et al., 2024)) simplifies training from a variational bound to a reweighted cross-entropy over masked positions
$\small{\mathcal{L}_{\text{MDM}}=\int_{0}^{1}\frac{1}{t}\mathbb{E}_{q_{t|0}(\mathbf{x}_{t}|\mathbf{x}_{0})}\left[\sum_{i:x_{t}^{i}=\text{MASK}}\!\!\!\!\!\!\!\!-\log p_{\theta}(x_{0}^{i}|\mathbf{x}_{t})\right]dt.}$ (2)
This formulation scales to LLMs as DLMs, with LLaDA (Nie et al., 2025) and Dream-7B (Ye et al., 2025) matching autoregressive performance while enabling parallel decoding and flexible infilling.
The analytical reverse of the forward process defined in Equation 1 is computationally inefficient for generation, as it typically modifies only a single token at each step (Campbell et al., 2022; Lou et al., 2023). A common strategy to accelerate sampling is to employ an MCMC-based approximation of the reverse process (Nie et al., 2025), enabling the model to update multiple masked tokens in a single step when transitioning from a noise level $t$ to an earlier level $s<t$. This yields the following factorized form:
$q_{s|t}=\prod_{i=1}^{L}q_{s|t}(\boldsymbol{x}_{s}^{i}\mid\boldsymbol{x}_{t}).$
$\begin{cases}\qquad\qquad 1&\text{if }\boldsymbol{x}_{s}^{i}=\boldsymbol{x}_{t}^{i}\neq[\text{MASK}],\\[4.0pt] \qquad\qquad\dfrac{s}{t}&\text{if }\boldsymbol{x}_{s}^{i}=\boldsymbol{x}_{t}^{i}=[\text{MASK}],\\[6.0pt] \dfrac{t-s}{t}\,q_{0|t}(\boldsymbol{x}_{s}^{i}\mid\boldsymbol{x}_{t})&\text{if }\boldsymbol{x}_{t}^{i}=[\text{MASK}]\neq\boldsymbol{x}_{s}^{i}.\end{cases}$ (3)
Here, $q_{0|t}(\boldsymbol{x}_{s}^{i}\mid\boldsymbol{x}_{t})$ denotes the model-provided distribution over the vocabulary for predicting a non-$\text{MASK}]$ token when $\boldsymbol{x}_{t}^{i}$ is masked. In conditional generation settings e.g., generating a response $\boldsymbol{x}_{0}$ given a prompt $p$, the reverse diffusion process in Equation 3 must be adapted. In this case, the model’s predictive distribution for unmasking a token becomes prompt-conditioned: $q_{0|t}(\boldsymbol{x}_{s}^{i}\mid\boldsymbol{x}_{t},p),$ reflecting that token predictions now depend on both the intermediate noised sequence and the conditioning prompt.
3.2 Diffusion Mamba Language Models
We propose DiffuMamba, an MDM whose denoiser replaces the Transformer encoder with a bidirectional state-space Mamba (BiMamba) backbone (Zhu et al., 2024; Sahoo et al., 2024; Schiff et al., 2024), preserving the probabilistic structure of masked discrete diffusion while enabling linear-time inference and substantially reduced memory overhead during multi-step denoising. We refer to transformer based DLM in our work as DiffuTran for simplicity.
Figure 1: Schematic diagram of our proposed DiffuMamba architecture where mixer blocks replaces attention layers with bidirectional Mamba layers. For DiffuMamba-H we have interleaved attention layers after every $N$ Mamba layers. Attention provides global token interactions while Mamba enables efficient state space sequence modeling, allowing the hybrid denoiser to capture both long-range dependencies and local temporal dynamics with significantly improved efficiency. In our experiments, we fix $N=5$.
Architecture details.
We now describe the BiMamba denoiser architecture used in DiffuMamba also shown in Figure 1. Let $\mathbf{x}\in\mathbb{R}^{B\times L\times d}$ denote a batch of token embeddings, where $B$ is the batch size, $L$ the sequence length, and $d$ the hidden dimension. Each block is composed of two independent Mamba layers: one processes the sequence in the forward direction, and the other processes the sequence in the reverse direction
$\mathbf{h}^{\rightarrow}_{i}$ $=A_{f}\ast\mathbf{h}^{\rightarrow}_{i-1}+B_{f}\ast\mathbf{x}_{i},\quad i=1,\dots,L,$ (4)
$\mathbf{h}^{\leftarrow}_{i}$ $=A_{b}\ast\mathbf{h}^{\leftarrow}_{i+1}+B_{b}\ast\mathbf{x}_{i},\quad i=L,\dots,1.$ (5)
Here $A_{f},B_{f},A_{b},B_{b}\in\mathbb{R}^{k\times d}$ are learnable state-transition kernels implemented as 1D causal and anti-causal convolutions or efficient scan operations (Gu & Goel, 2023). The two directional feature streams are fused through simple additive integration
$\text{Mamba}(x_{i})=\mathbf{h}_{i}=\mathbf{h}^{\rightarrow}_{i}+\mathbf{h}^{\leftarrow}_{i},$ (6)
providing a symmetric context representation while maintaining numerical stability. The resulting hidden sequence $\mathbf{H}=(\mathbf{h}_{1},\dots,\mathbf{h}_{L})$ is then normalized and passed through a lightweight feed-forward projection before residual addition.
Each diffusion block applies noise-conditioned Mamba mixing followed by an MLP refinement with residual connections, and the resulting hidden states are propagated through a stack of such blocks to predict token logits at each denoising step. The model defines a categorical distribution over clean tokens and is trained using the standard masked diffusion objective, preserving probabilistic consistency with absorbing-state discrete diffusion. We defer the full architectural formulation, conditioning mechanism, and training details to Appendix A.1. By replacing attention with bidirectional state-space dynamics, DiffuMamba achieves efficient and expressive denoising with linear scaling in both sequence length and memory.
Recent results in large-scale autoregressive modeling show that hybrid architectures (Lieber et al., 2024; Bae et al., 2025; Zuo et al., 2025; Wang et al., 2025a) combining attention and state-space layers can outperform both pure Transformer and pure Mamba models under similar compute budgets, leveraging complementary strengths in local recurrence and global dependencies. Motivated by these findings, we investigate hybrid DLMs by interleaving bidirectional Mamba and Transformer blocks, inserting one attention mixer every $N$ Mamba mixer blocks [^1] NN is a hyperparameter that can be tuned to balance model performance and inference throughput. (with $N=5$ in all experiments). Extending this paradigm from AR to diffusion, we examine whether similar complementarities hold under masked denoising objectives.
Table 1: Zeroshot Perplexities (PPL $\downarrow$) across Benchmarks for Different Model Sizes and Configurations. Best PPL are highlighted in blue and second best are underlined. We observe that DiffuMamba-H (1.3B) achieves the best overall performance across all configurations, with DiffuMamba as the next strongest model. At smaller scales (240M), however, DiffuMamba delivers performance comparable to its attention-based diffusion counterpart, indicating that the advantages of hybridization become more pronounced at larger model sizes.
Model Size Configuration PTB WikiText LM1B Lambada AG News PubMed ArXiv Avg.
240M DiffuTran 153.47 49.06 142.25 49.06 105.06 29.83 43.79 81.79
DiffuMamba 99.04 51.35 169.92 46.66 112.59 33.98 33.20 78.11
DiffuMamba-H 147.97 50.96 145.25 46.43 114.20 31.94 30.99 81.11
0.5B DiffuTran 112.69 40.76 103.51 39.73 81.49 25.51 25.37 61.15
DiffuMamba 116.29 42.56 117.42 40.31 90.10 27.60 27.79 65.87
DiffuMamba-H 110.42 40.25 100.59 39.30 78.38 26.12 25.84 60.13
1.3B DiffuTran 101.14 36.63 99.72 36.73 67.53 23.03 23.25 55.29
DiffuMamba 99.04 34.74 102.01 36.04 68.75 22.35 22.98 55.13
DiffuMamba-H 96.51 31.92 92.83 34.04 63.22 19.05 20.67 51.18
4 Experiments
We evaluate three architectures under an identical diffusion objective, masking process, and noise schedule to ensure a controlled comparison. The first configuration, DiffuTran, employs an MHA-based mixer in every diffusion block; the second, DiffuMamba, replaces all such mixers with bidirectional Mamba blocks; and the third, DiffuMamba-H, adopts a hybrid design that interleaves MHA and Mamba blocks. Section 3.2 details the corresponding design principles. Across all variants, the overall architecture, data flow, and conditioning mechanisms remain identical, with only the internal mixer type varying, as illustrated in Figure 1. To maintain comparable parameter budgets, the MLP expansion ratio in DiffuMamba and hybrid configurations is set to half that of the MHA-only model, while other hyperparameters, such as the number of blocks, hidden size, etc are kept constant as detailed in Table 5. This design ensures a fair, architecture-isolated evaluation of modeling quality, efficiency, and inference throughput. Our goals are to answer:
  1. 1.
    Performance & Scaling: Can Mamba-based DLMs reach or even surpass the performance of standard Transformer DLMs at different parameter scales and training tokens scales?
  2. 2.
    Throughput and Latency: Does an SSM backbone yield higher tokens/sec and lower per-step latency with identical inference settings?
4.1 Data, Tokenization, and Pretraining Setup
We evaluate DiffuMamba and DiffuMamba-H models as generative LMs by comparing their performance against DiffuTran at three parameter scales: 240M, 0.5B, and 1.3B. We further compare throughput and wall-clock model latency at the largest 1.3B scale. All models are trained on the DCLM dataset (Li et al., 2024) using the GPT2 tokenizer (Radford et al., 2019), with a fixed context length of 1,024 tokens. We train all models with a fixed log-linear noise schedule and the masked diffusion objective, under approximate Quokka optimal compute budget (Ni et al., 2025b) as given in the Appendix, Table 5 and Table 6. For measuring inference throughput, we record the wall-clock model latency and tokens-per-second throughput on a single NVIDIA H100 GPU using bf16 precision and PyTorch backend using CUDA-graphs. Each measurement averages over 5 full diffusion decoding runs following 3 warm-up iterations.
4.2 Modeling Quality
[…]
Config 240M 0.5B 1.3B
Chinchilla Quokka Chinchilla Quokka Chinchilla Quokka
DiffuTran 43.82 32.11 31.03 25.07 25.01 22.72
DiffuMamba 44.09 33.01 30.78 25.46 23.96 21.41
DiffuMamba-H 42.67 32.49 29.06 25.14 22.89 20.17
We first measure validation perplexity (Val. PPL) for different configuration of models as shown in Table 2. Across all model sizes, hybrid DiffuMamba-H consistently outperform their pure-attention counterparts under both Chinchilla and Quokka compute budgets. At the 1.3B scale, DiffuTran attains 25.01 and 22.72 PPL, but DiffuMamba-H achieves the best results with 22.89 and 20.17, delivering roughly a 2% perplexity reduction. At smaller scales (240M and 0.5B), DiffuMamba-H remain competitive, often securing the best or second-best perplexities, demonstrating that hybridization provides consistent gains across pre-training regimes.
Table 3: Asymptotic inference efficiency at batch size B=1B{=}1. We report FLOPs ($F$), memory operations ($M$), arithmetic intensity (AI=F/M=F/M), and throughput ($T$). $L$ is sequence length, $d$ hidden dimension, $K$ diffusion steps, AA Mamba state size, $G$ block size, and $p$ the step scaling factor. Cost includes block diffusion steps and cache recomputation.
Model FLOPs ($F$) Memory ($M$) AI ($F/M$) Throughput ($T$)
AR $\mathcal{O}(Ld^{2}+L^{2}d)$ $\mathcal{O}(Ld^{2}+L^{2}d)$ $\mathcal{O}(1)$ $\mathcal{O}\!\left(\dfrac{1}{Ld+d^{2}}\right)$
DiffuTran $\mathcal{O}(KLd^{2}+KL^{2}d)$ $\mathcal{O}(KLd+Kd^{2})$ $\mathcal{O}(L)$ $\mathcal{O}\!\left(\dfrac{C_{\max}}{KLd+Kd^{2}}\right)$
DiffuMamba $\mathcal{O}(KLd^{2}+KLdA)$ $\mathcal{O}(KLd+Kd^{2}+KLA)$ $\mathcal{O}\!\left(\dfrac{dL}{d+L}\right)$ $\mathcal{O}\!\left(\dfrac{L}{KLd+Kd^{2}}\right)$
DiffuMamba-H $\mathcal{O}(KLd^{2}+KL^{2}d)$ $\mathcal{O}(KLd+Kd^{2})$ $\mathcal{O}(L)$ $\mathcal{O}\!\left(\dfrac{C_{\max}}{KLd+Kd^{2}}\right)$
DiffuMamba-H + Fast-dLLM $\mathcal{O}\!\left(KGd^{2}+KGLd+\dfrac{L}{G}(Ld^{2}+L^{2}d)\right)^{\dagger}$ $\mathcal{O}(KLd+Kd^{2})$ $\mathcal{O}\!\left(\dfrac{pL}{G}\right)$ $\mathcal{O}\!\left(\dfrac{C_{\max}}{KLd+Kd^{2}}\right)$
We next assess models’ ability to generalize zero-shot to held-out datasets. The results are presented in Table 1. Following Sahoo et al. (2024), we evaluate PPL upper bounds on Penn Treebank (PTB; Marcus et al., 1993), WikiText (Merity et al., 2017), LM1B (Chelba et al., 2013), Lambada (Paperno et al., 2016), AG News (Zhang et al., 2015), and Pubmed/Arxiv subsets of Scientific Papers (Cohan et al., 2018). All models are evaluated without fine-tuning.
At the 240M scale, DiffuTran outperforms both DiffuMamba and DiffuMamba-H on 4 out of 7 datasets, indicating that at smaller scales, Mamba-based models struggle to generalize effectively. This trend can also be observed in Table 2, where at small scale DiffuTran has higher Val. PPL. At 0.5B, the benefits of hybridization become immediately evident: DiffuMamba-H outperforms all other variants on 5 out of 7 datasets, establishing a clear performance lead. The advantage becomes even more pronounced at 1.3B, where DiffuMamba-H surpasses DiffuTran across all datasets, with DiffuMamba ranking second in most cases (5 out of 7) indicating that Mamba’s sequence-modeling inductive bias scales more effectively than pure attention for diffusion denoising.
These results highlight a key trend: while Mamba enables efficient long-range token mixing, adding interleaved attention layers in DiffuMamba-H captures complementary global dependencies consistent with findings in AR linear attention (Wang et al., 2025a). This hybrid design consistently improves generalization at larger scales.
We also report downstream eval results for 1.3B models in Table 4. As expected, the absolute scores remain modest at this scale, reflecting the difficulty of these reasoning and knowledge-intensive for base models at 1.3B scale. Nonetheless, a consistent trend emerges across all tasks: DiffuMamba and DiffuMamba-H clearly outperform DiffuTran by $\approx\boldsymbol{4\%}$ on average, indicating that linear-time state space modeling provides a stronger denoising backbone for diffusion-based LMs. Further hybridizing with attention layers, as in DiffuMamba-H  produces the strongest results, suggesting that a small degree of explicit cross-token interaction complements the Mamba backbone. Even at low model capacity where downstream metrics are weak, Mamba based DLM clearly outperform attention based DiffuTran, demonstrating the effectiveness of state space driven denoising and its superior inference throughput.
Takeaway. From 240M to 1.3B models Mamba-based DLMs match or improve upon MHA-based ones in validation and zero-shot PPL, with hybrid DiffuMamba-H yielding the most consistent gains at larger scale.
Table 4: Zero-shot downstream accuracy ($\uparrow$) on reasoning and commonsense benchmarks for 1.3B models. At this scale, overall performance is modest, as 1.3B models struggle on these challenging benchmarks. Both DiffuMamba and DiffuMamba-H consistently outperform the attention-based diffusion baseline (DiffuTran), with DiffuMamba-H achieving the strongest results. Best results are shown in blue; second-best are underlined.
Config (1.3B) OBQA HS PIQA LOQA ARC Avg.
DiffuTran 26.61 34.97 60.64 21.86 24.98 33.81
DiffuMamba 32.12 37.74 62.56 29.87 28.33 38.12
DiffuMamba-H 33.87 38.02 59.13 32.35 27.83 38.24
[…]
4.3 Inference Throughput
Figure 3: Latency decomposition per denoising step for a single forward pass across different model configurations, obtained by fitting a non-negative quadratic model $t(L)=aL^{2}+bL+c$ to the measured latency curves. The quadratic coefficient $a$ captures attention-like pairwise interactions, while the linear coefficient $b$ reflects token-wise computations (e.g., MLP/SSM), and $c$ denotes constant overheads. For intermediate sequence lengths, linear and constant terms dominate across models. At large lengths, the superior scaling of DiffuMamba stems from its negligible quadratic component, whereas DiffuTran becomes dominated by the quadratic term. Although DiffuMamba-H reduces latency relative to DiffuTran, its scaling remains governed by attention.
LABEL:fig:inf:a and LABEL:fig:inf:b compare inference throughput between the attention-based diffusion model, DiffuTran and our proposed DiffuMamba and DiffuMamba-H as sequence length ($L$) increases from $64$ to $65536$. Following Nie et al. (2025); Wu et al. (2025), we fix batch size $B=1$. We are interested in evaluating inference efficiency under different decoding paradigms, hence we include Fast-dLLM (Wu et al., 2025) – a training free block diffusion decoding paradigm that relies on KV-cache for inter-block diffusion and recomputes KV-cache after each generation block. Throughput ($T$) is measured in tokens per second as the number of generated tokens divided by total wall-clock decoding time. Unlike prior works that use a fixed number of denoising steps (e.g., $K=128$ in Nie et al. (2025)), we scale the number of steps with the sequence length as $K=L/p$ to enable long-context evaluation, and report results for $p\in\{8,16\}$. For Fast-dLLM we fix the block size to $G=32$ following Wu et al. (2025), yielding denoising steps per block, $k=32/p$.
It can be observed that for moderate sequence lengths ($L\leq 2\mathrm{K}$), DiffuTran+Fast-dLLM achieves the highest throughput among all models. In this regime, DiffuMamba and DiffuTran exhibit competitive and nearly identical performance. As the sequence length increases, throughput gradually decreases for all methods, consistent with the asymptotic analysis in Table 3. In particular, throughput becomes dominated by the hidden dimension $d$ rather than by $L$, staying approximately constant in this regime.
Beyond $2K$ tokens in LABEL:fig:inf:a and LABEL:fig:inf:b, the throughput of DiffuTran degrades sharply, while DiffuMamba (memory bandwidth bound (Baruah et al., 2025)) and DiffuMamba-H (reduces FLOPs per forward) experience a substantially slower decline. This trend aligns with the asymptotic throughput analysis in Table 3, which predicts more favorable scaling behavior for DiffuMamba and DiffuMamba-H. When L>d(1920)L>d\;(1920) (approximately at $L=2048$), the diffusion model enters a compute-saturated regime. In this regime, the FLOPs per token grow with $L$, while the sustained FLOPs/s remain bounded by the peak hardware capacity $C_{\max}$, resulting in a decline in throughput. Beyond this point, DiffuTran exhibits the steepest throughput collapse due to the combined effects of quadratic attention cost and linearly increasing denoising steps, resulting in $T=\mathcal{O}(\frac{C_{max}}{KLd+Kd^{2}})\approx\mathcal{O}(\frac{1}{L^{2}})$, where $K=L/p$, scaling at long context lengths. In contrast, DiffuMamba is memory bound. As a result, for $L>d$, throughput scales as $\mathcal{O}(\frac{L}{Kd^{2}+KLd})$, where $K=L/p$, thus $T=\mathcal{O}(\frac{p}{d^{2}+Ld})=O(\frac{1}{L})$. This gives slower throughput degradation, retaining a $\boldsymbol{8.2\times}$ and $\boldsymbol{4.3\times}$ advantage at $K=L/16$ denoising steps for DiffuMamba and DiffuMamba-H respectively over DiffuTran at 6565K tokens.
In LABEL:fig:infb:a and LABEL:fig:infb:b, we evaluate all DLMs using a block-autoregressive inference that reuses cached representationsacross successive generation blocks (Wang et al., 2025b; Arriola et al., 2025). Concretely, once a block of tokens is denoised, the corresponding cache is retained and directly reused when processing the next block, avoiding repeated forward passes over previously generated tokens. As generation proceeds, caches accumulate incrementally across blocks, enabling efficient long-context inference.
[…]
To further understand the throughput trends in Figure 2, we analyze the per-forward-pass latency as a function of sequence length $L$. From each benchmark record, we compute the mean latency per denoising step as $t(L)\;=\;\frac{\texttt{Model Total Time}}{\texttt{steps}}$ which isolates the cost of a single model forward pass at sequence length $L$. We model the measured latency using a non-negative quadratic decomposition, $t(L)\;=\;aL^{2}+bL+c$ s.t. $a,b,c\geq 0,$ where the quadratic term $aL^{2}$ captures attention-like pairwise token interactions, the linear term $bL$ captures token-wise computations (e.g., MLP/SSM/projections), and the constant term $c$ captures sequence-length–independent overheads.
The fitted equations make the scaling behavior explicit as demonstrated in Figure 3. Attention-based diffusion (DiffuTran) has a comparatively larger quadratic coefficient, so the quadratic component $aL^{2}$ becomes dominant as $L$ grows. In contrast, Mamba-based diffusion (DiffuMamba) exhibits a much smaller quadratic coefficient; over the evaluated range, latency is largely explained by the linear and constant components. Hybrid interleavings (DiffuMamba-H variants) sit between these regimes, retaining a controlled quadratic contribution due to periodic attention insertion.
Taken together, this latency decomposition clarifies the fundamental source of the throughput advantages observed in Figure 2. By isolating the role of quadratic attention costs, our analysis explains why Mamba-based diffusion exhibits more favorable scaling and why hybrid designs interpolate smoothly between the two regimes. These insights motivate the architectural choices we summarize next and provide a principled foundation for the conclusions that follow.
Takeaway Mamba-based DLMs achieve higher throughput than MHA-based ones under all inference algorithms. Block caching removes quadratic recomputation in Transformers pushing DLMs’ throughput far beyond AR. Mamba-based block-AR DLMs with constant memory requirement achieve best throughput in this memory-bound setting.
5 Conclusion
This work demonstrates the feasibility of DLMs built with Mamba backbones. We introduce DiffuMamba, the first diffusion LM that relies exclusively on linear state-space mixers instead of multi-head attention, and DiffuMamba-H, a hybrid variant that interleaves Mamba-2 and attention layers to capture complementary global context. Across model scales from 240M to 1.3B parameters, both approaches match or surpass the Transformer-based baseline DiffuTran in language modeling performance.
Our throughput analysis demonstrates that Mamba-based DLMs scale more favorably than MHA-based DLMs across a range of inference algorithms. In particular, DLMs relying on block cache reuse yield the strongest performance across the entire context length spectrum efficiently combining the multi-token generation advantages of DLMs and preventing quadratic scaling by fixing the representations of the past tokens (e.g. using KV-cache). While DiffuMamba + Block Cache emerges as the most efficient design for long sequence generation, achieving optimal performance in this regime requires training with block-diffusion inductive biases or incorporating distillation objectives as in Wang et al. (2025b). As the primary goal of this work is to establish feasibility and characterize the efficiency advantages of DLMs with linear mixers, we leave the training of block-cached hybrid models at useful scales and the exploration of alternative linear mixers to future work. We summarize the main limitations of this work in Appendix B.
Acknowledgment
We would also like to acknowledge funding from the MITACS Accelerate and NSERC Discovery. We also acknowledge resources provided by Compute Canada, Calcul Québec, and Mila.
Appendix A Model Configurations
Model configurations for DiffuTran, DiffuMamba and DiffuMamba-H in our experiments. All variants are trained under the same diffusion objective, noise schedule, and tokenization. The only difference lies in the internal mixer architecture. The MLP expansion ratio in DiffuMamba and DiffuMamba-H configurations is reduced by half for comparable parameter budgets. In the hybrid architecture, an attention layer is inserted every $N$ layers; we set $N=5$ in all experiments. In Table 6, we list down all the training and inference hyperparamters.
Table 5: Model configurations for DiffuTran, DiffuMamba and DiffuMamba-H.
Model Params (B) Layers $\mathbf{d_{\text{model}}}$ $\mathbf{d_{\text{mlp}}}$ $\mathbf{d_{\text{head}}}$ $\mathbf{d_{\text{state}}}$ Context Len. #\# Tokens
DiffuTran-0.24B 0.24 24 960 960*4 32 - 1024 25B
DiffuMamba-0.29B 0.29 24 960 960*2 32 128 1024 25B
DiffuMamba-H-0.26B (K=5) 0.26 24 960 960*2 32 128 1024 25B
DiffuTran-0.5B 0.52 36 960 960*4 32 - 1024 50B
DiffuMamba-0.6B 0.67 36 960 960*2 32 128 1024 50B
DiffuMamba-H-0.6B (K=5) 0.61 36 960 960*2 32 128 1024 50B
DiffuTran-1.3B 1.3 24 1920 1920*4 32 - 1024 120B
DiffuMamba-1.6B 1.6 24 1920 1920*2 32 128 1024 120B
DiffuMamba-H-1.5B (K=5) 1.5 24 1920 1920*2 32 128 1024 120B
A.1 Architecture Details of DiffuMamba
We present here the architectural details of DiffuMamba  an MDM whose denoiser is built on a bidirectional state-space Mamba (BiMamba) backbone. DiffuMamba preserves the probabilistic structure of masked discrete diffusion while replacing the Transformer encoder with a bidirectional Mamba mixer, enabling linear-time inference and substantially lower memory overhead during multi-step denoising.
Let $\mathbf{x}\in\mathbb{R}^{B\times L\times d}$ denote a batch of token embeddings, where $B$ is the batch size, $L$ the sequence length, and $d$ the hidden dimension. Each block is composed of two independent Mamba layers: one processes the sequence in the forward direction, and the other processes the sequence in the reverse direction
$\mathbf{h}^{\rightarrow}_{i}$ $=A_{f}\ast\mathbf{h}^{\rightarrow}_{i-1}+B_{f}\ast\mathbf{x}_{i},\quad i=1,\dots,L,$ (7)
$\mathbf{h}^{\leftarrow}_{i}$ $=A_{b}\ast\mathbf{h}^{\leftarrow}_{i+1}+B_{b}\ast\mathbf{x}_{i},\quad i=L,\dots,1.$ (8)
Here $A_{f},B_{f},A_{b},B_{b}\in\mathbb{R}^{k\times d}$ are learnable state-transition kernels implemented as 1D causal and anti-causal convolutions or efficient scan operations (Gu & Goel, 2023). The two directional feature streams are fused through simple additive integration
$\text{Mamba}(x_{i})=\mathbf{h}_{i}=\mathbf{h}^{\rightarrow}_{i}+\mathbf{h}^{\leftarrow}_{i},$ (9)
providing a symmetric context representation while maintaining numerical stability. The resulting hidden sequence $\mathbf{H}=(\mathbf{h}_{1},\dots,\mathbf{h}_{L})$ is then normalized and passed through a lightweight feed-forward projection before residual addition.
Each Diffusion Block employs timestep-conditioned adaptive layer normalization (AdaLN) to inject the diffusion noise level into the hidden activations. A small MLP maps the scalar timestep $t$ to a continuous embedding $\tau_{t}=\text{MLP}(t)\in\mathbb{R}^{d_{c}}$, which modulates both the mixer and MLP sublayers:
$\text{AdaLN}(\mathbf{x};\tau_{t})=\gamma_{t}\cdot\frac{\mathbf{x}-\mu(\mathbf{x})}{\sigma(\mathbf{x})}+\beta_{t},\quad(\gamma_{t},\beta_{t})=W_{\text{cond}}\tau_{t}.$ (10)
This conditioning allows the denoiser to adapt its internal recurrence dynamics to varying noise levels across diffusion steps.
Each block then applies Mamba mixing followed by an MLP refinement as shown in Figure 1:
$\mathbf{y}_{\text{mixer}}$ $=\text{Mamba}(\text{AdaLN}(\mathbf{x};\tau_{t}))+\mathbf{x},$ (11)
$\mathbf{y}_{\text{mlp}}$ $=\text{MLP}(\text{AdaLN}(\mathbf{y}_{\text{mixer}};\tau_{t}))+\mathbf{y}_{\text{mixer}}.$ (12)
The final output $\mathbf{y}_{\text{mlp}}$ is passed to the next block ($\mathrm{B}$) in the stack. During denoising, given masked input embeddings $\mathbf{x}_{t}$ and timestep embedding $\tau_{t}$, the model predicts token logits via the full diffusion stack:
$\mathbf{z}=f_{\theta}(\mathbf{x}_{t},\tau_{t})=\text{OutputProj}\!\left(\mathrm{B}_{N}(\cdots\mathrm{B}_{1}(E(\mathbf{x}_{t}),\tau_{t})\cdots)\right).$ (13)
The conditional distribution over clean tokens is
$p_{\theta}(\mathbf{x}_{0}|\mathbf{x}_{t},t)=\prod_{i=1}^{L}\text{Cat}(x_{0}^{i};\text{softmax}(z_{i})).$ (14)
Training minimizes the masked diffusion objective (Eq. 2), preserving probabilistic consistency with absorbing-state discrete diffusion.
Table 6: Key training and inference hyperparameters used across all experiments unless stated otherwise.
Category Setting
Backbone Transformer (DiffuTran), Mamba (DiffuMamba), Hybrid (DiffuMamba-H)
Diffusion Type Absorbing-state masked diffusion
Parameterization Substitution (subs)
Noise Conditioning Log Linear
Global Batch Size 512
Precision bf16
EMA Decay 0.9999
Antithetic Sampling Enabled
Sampling $\epsilon$ $10^{-3}$
Gradient Clipping 1.0
Optimizer Adam
Max Learning Rate $1\times 10^{-4}$
Min Learning Rate $1\times 10^{-6}$
Weight Decay 0.1
Adam Betas (0.9, 0.95)
Adam $\epsilon$ $10^{-8}$
LR scheduler Cosine
Denoising Steps factor (p) {8, 16}
Block Length 32
Inference Batch Size 1
hidden dimension 960
#\# blocks 24
#\# heads 24
bidirectional weight tie False
mamba state dimension 128
mamba conv dimension 4
expansion factor 2
interleaving attention (N) 5
Appendix B Limitations
While our results demonstrate strong inference gains, the joint training of diffusion language models with Mamba backbones and block cache reuse remains unexplored. Our evaluation therefore isolates the effect of backbone choice and inference algorithm on throughput and latency, but does not reflect potential gains from training models explicitly for the fastest decoding regimes. In particular, training models explicitly optimized for the fastest inference algorithms such as block diffusion with cache reuse alongside the exploration of alternative linear-time token mixers, remains a natural and immediate next step.
Further our throughput measurements assume fixed acceptance rates through the parameter $p$, which controls the number of denoising steps $K=L/p$. Although diffusion models achieve their best quality when $p=1$, this setting effectively reduces diffusion decoding to autoregressive-style computation and defeats the primary motivation of diffusion-based generation. Our choice of moderate $p$ values reflects the intended efficiency regime of DLMs, but alternative acceptance behaviors may lead to different trade-offs.
Finally, all experiments are conducted at batch size $B=1$. This setting is consistent with prior work (Nie et al., 2025; Wu et al., 2025) on long-context decoding and reflects the practical constraints of GPU memory at extreme sequence lengths. Evaluating larger batch sizes in the long-context regime remains an important direction for future study.